OpenStudy (anonymous):

Two tanks contain equal amounts of water. They are connected by a pipe and 3000 litres of water is pumped from one tank to the other. One tank contains 6 times as much water as the other. How many litres of water did each tank contain originally?

OpenStudy (anonymous):

Call the tanks A and B. Originally A=B. Say the water was pumped from A to B. That means subtract 3000 liters from A (A-3000) and add 3000 liters to B (B+3000). B+3000 is 6 times A-3000. 6(A-3000) = B+3000 ---> 6A - 18000 = B + 3000. Remember, that A = B. So substitute B into the equation for A. 6B - 18000 = B + 3000 5B = 21000 B = 21000/5 = 4200 = A. Both tanks originally held 4200 liters.